### v2

# Weyl groups and rigidity of von Neumann algebras

Abstract.
Let $G$ be a noncompact semisimple algebraic group with trivial center, $S<G$ a maximal split torus, $H<G$ the centralizer of $S$ in $G$ and $\Gamma<G$ an irreducible lattice. Consider the group measure space von Neumann algebra $\mathscr{M}=\operatorname{L}(\Gamma\curvearrowright G/H)$ associated with the nonsingular action $\Gamma\curvearrowright G/H$ and regard the group von Neumann algebra $M=\operatorname{L}(\Gamma)$ as a von Neumann subalgebra $M\subset\mathscr{M}$. We show that the group $\operatorname{Aut}_{M}(\mathscr{M})$ of all unital normal $\ast$-automorphisms of $\mathscr{M}$ acting identically on $M$ is isomorphic to the Weyl group $\mathscr{W}_{G}$ of the semisimple algebraic group $G$. Our main theorem is a noncommutative analogue of a rigidity result of Bader–Furman–Gorodnik–Weiss for group actions on algebraic homogeneous spaces and moreover gives new insight towards Connes’ rigidity conjecture for higher rank lattices.
Key words and phrases:
Algebraic groups; Irreducible lattices; von Neumann algebras; Weyl group
2020 Mathematics Subject Classification:
20G25, 22D25, 37A40, 46L10, 46L55
CH is supported by ERC Advanced Grant NET 101141693
AI is supported by NSF grants DMS-2153805 and DMS-2451697.
1. Introduction and statement of the main results
In order to state our main results, we use the following terminology regarding algebraic groups. We refer to [Bo91, Ma91] for further details.
Terminology.
Let $k$ be a local field of characteristic zero, that is, $k$ is equal to $\mathbb{R}$, $\mathbb{C}$ or a finite extension of $\mathbb{Q}_{p}$ for some prime number $p$. Let $\mathbf{G}$ be a Zariski connected simply connected $k$-isotropic almost $k$-simple algebraic $k$-group, $\mathbf{S}<\mathbf{G}$ a maximal $k$-split torus and $\mathbf{P}<\mathbf{G}$ a minimal parabolic $k$-subgroup such that $\mathbf{H}=\mathscr{Z}_{\mathbf{G}}(\mathbf{S})<\mathbf{P}$. The dimension of $\mathbf{S}$ is called the $k$-rank of $\mathbf{G}$ and is denoted by $\operatorname{rk}_{k}(\mathbf{G})$. Since $\mathbf{G}$ is $k$-isotropic, we have $\operatorname{rk}_{k}(\mathbf{G})\geq 1$. The centralizing $k$-subgroup $\mathscr{Z}_{\mathbf{G}}(\mathbf{S})$ is the Zariski connected component of the normalizing $k$-subgroup $\mathscr{N}_{\mathbf{G}}(\mathbf{S})$. The finite group $\mathscr{W}_{\mathbf{G}}=\mathscr{N}_{\mathbf{G}}(\mathbf{S})/\mathscr{Z}_{\mathbf{G}}(\mathbf{S})$ is called the Weyl group of $\mathbf{G}$ relative to $k$. Every coset of $\mathscr{N}_{\mathbf{G}}(\mathbf{S})/\mathscr{Z}_{\mathbf{G}}(\mathbf{S})$ is represented by an element rational over $k$, that is, $\mathscr{N}_{\mathbf{G}}(\mathbf{S})=\mathscr{N}_{\mathbf{G}}(\mathbf{S})(k)\cdot\mathscr{Z}_{\mathbf{G}}(\mathbf{S})$. In the case when $\mathbf{G}=\operatorname{SL}_{n}$ for $n\geq 2$, we may assume that $\mathbf{S}<\mathbf{G}$ is the group of diagonal matrices, we have $\mathbf{H}=\mathbf{S}$ and $\mathscr{W}_{\mathbf{G}}\cong\mathfrak{S}_{n}$.
We introduce the following notation which we will be using throughout our paper. Let $d\geq 1$. For every $i\in\{1,\dots,d\}$, let $k_{i}$ be a local field of characteristic zero, $\mathbf{G}_{i}$ a Zariski connected simply connected $k_{i}$-isotropic almost $k_{i}$-simple algebraic $k_{i}$-group as above and choose $\mathbf{S}_{i}<\mathbf{P}_{i}<\mathbf{G}_{i}$ accordingly so that $\mathbf{H}_{i}=\mathscr{Z}_{\mathbf{G}_{i}}(\mathbf{S}_{i})<\mathbf{P}_{i}$. Set $G=\prod_{i=1}^{d}\mathbf{G}_{i}(k_{i})$, $S=\prod_{i=1}^{d}\mathbf{S}_{i}(k_{i})$, $H=\prod_{i=1}^{d}\mathbf{H}_{i}(k_{i})$, $P=\prod_{i=1}^{d}\mathbf{P}_{i}(k_{i})$ and $\mathscr{W}_{G}=\prod_{i=1}^{d}\mathscr{W}_{\mathbf{G}_{i}}$. Endow the homogeneous space $G/H$ (resp. $G/P$) with its unique $G$-invariant measure class. Since $G$ and $H$ are both unimodular, the homogeneous space $G/H$ carries a $G$-invariant $\sigma$-finite infinite measure $m$ whose measure class coincides with the unique $G$-invariant measure class on $G/H$. It is well known that as nonsingular $G$-spaces, we have a natural isomorphism $G/H\cong G/P\times G/P$, where the composition with the projection onto the first factor corresponds to the factor map $G/H\to G/P$. Let $\Gamma<G$ be an irreducible lattice and set $\Lambda=\Gamma/\mathscr{Z}(\Gamma)$. Since $\mathscr{Z}(\Gamma)<H\cap\mathscr{Z}(G)$, we have $\mathscr{Z}(\Gamma)<\ker(\Gamma\curvearrowright G/H)$ and so we may consider the well-defined nonsingular action $\Lambda\curvearrowright G/H$. Using the terminology on algebraic groups as above and since $H<G$ is noncompact and $\Gamma<G$ is irreducible, Moore’s ergodicity theorem implies that $\Lambda\curvearrowright G/H$ is ergodic (see [HM77]). Moreover, Lemma 2.3 below implies that $\Lambda\curvearrowright G/H$ is essentially free.
As a consequence of the main results of [BFGW12] (see Theorem 2.2 below for further details), any $\Lambda$-equivariant nonsingular automorphism of $G/H$ is automatically $G$-equivariant and so we have the following isomorphisms
$\operatorname{Aut}_{\Lambda}(G/H)=\operatorname{Aut}_{\Gamma}(G/H)=\operatorname{Aut}_{G}(G/H)\cong\mathscr{W}_{G}.$
The aim of our paper is to prove a noncommutative analogue of the above isomorphism result. Before stating our main theorem, we need to introduce some further notation.
We denote by $\sigma:\Lambda\curvearrowright\operatorname{L}^{\infty}(G/H)$ the von Neumann algebraic action corresponding to the nonsingular action $\Lambda\curvearrowright G/H$. Denote by $M=\operatorname{L}(\Lambda)=\left\{u_{\gamma}\mid\gamma\in\Lambda\right\}^{\prime\prime}$ the group von Neumann algebra. Denote by $\mathscr{M}=\operatorname{L}(\Lambda\curvearrowright G/H)$ the group measure space von Neumann algebra associated with the nonsingular action $\Lambda\curvearrowright G/H$ together with its Cartan subalgebra $\mathscr{A}=\operatorname{L}^{\infty}(G/H)$ and regard $M\subset\mathscr{M}$ as a von Neumann subalgebra. For every $w=(w_{i})_{i}\in\mathscr{W}_{G}$, choose $n=(n_{i})_{i}\in\prod_{i=1}^{d}\mathscr{N}_{\mathbf{G}_{i}}(\mathbf{S}_{i})(k_{i})$ so that $w=(n_{i}\mathbf{H}_{i})_{i}\in\mathscr{W}_{G}$ and define $\theta_{w}\in\operatorname{Aut}(\mathscr{A})$ as the unique unital normal $\ast$-automorphism that satisfies $\theta_{w}(F)(gH)=F(gHn^{-1})$ for every $F\in\mathscr{A}$ and almost every $gH\in G/H$. Then we have $\theta_{w}\circ\sigma_{\gamma}=\sigma_{\gamma}\circ\theta_{w}$ for every $\gamma\in\Lambda$ and every $w\in\mathscr{W}_{G}$.
Denote by $\operatorname{Aut}_{M}(\mathscr{M})$ the group of all unital normal $\ast$-automorphisms $\Theta\in\operatorname{Aut}(\mathscr{M})$ such that $\Theta|_{M}=\operatorname{id}_{M}$. For every $w\in\mathscr{W}_{G}$, define $\Theta_{w}\in\operatorname{Aut}_{M}(\mathscr{M})$ as the unique unital normal $\ast$-automorphism that satisfies $\Theta_{w}|_{M}=\operatorname{id}_{M}$ and $\Theta_{w}|_{\mathscr{A}}=\theta_{w}$. Then $\rho:\mathscr{W}_{G}\to\operatorname{Aut}_{M}(\mathscr{M}):w\mapsto\Theta_{w}$ is an injective group homomorphism.
Our main result is the following novel rigidity phenomenon for the inclusion of von Neumann algebras $M=\operatorname{L}(\Lambda)\subset\operatorname{L}(\Lambda\curvearrowright G/H)=\mathscr{M}$ associated with the nonsingular action $\Lambda\curvearrowright G/H$.
Theorem A.
The group homomorphism $\rho:\mathscr{W}_{G}\to\operatorname{Aut}_{M}(\mathscr{M})$ is an isomorphism. In particular, we have
$\mathscr{W}_{G}\cong\operatorname{Aut}_{M}(\mathscr{M}).$
Since the inclusion $M\subset\mathscr{M}$ is irreducible (see Proposition 2.7 below), we have $\operatorname{Inn}_{M}(\mathscr{M})=\left\{\operatorname{Ad}(u)\mid u\in\mathscr{U}(M^{\prime}\cap\mathscr{M})\right\}=\{\operatorname{id}_{\mathscr{M}}\}$. This further implies that $\mathscr{W}_{G}\cong\operatorname{Aut}_{M}(\mathscr{M})=\operatorname{Out}_{M}(\mathscr{M})$. The striking feature of Theorem A is that even though the inclusion $M=\operatorname{L}(\Lambda)\subset\operatorname{L}(\Lambda\curvearrowright G/H)=\mathscr{M}$ tends to forget the nonsingular action $\Lambda\curvearrowright G/H$, it still retains the Weyl group $\mathscr{W}_{G}$ via the group isomorphism $\mathscr{W}_{G}\cong\operatorname{Aut}_{M}(\mathscr{M})$.
This result is related to Boutonnet–Houdayer’s noncommutative factor theorem [BH22] (see also [Ho21]). In that respect, set $\mathscr{B}=\operatorname{L}(\Lambda\curvearrowright G/P)$ and naturally regard $M\subset\mathscr{B}\subset\mathscr{M}$. Using the same setting as above and under the higher rank assumption $\sum_{i=1}^{d}\operatorname{rk}_{k_{i}}(\mathbf{G}_{i})\geq 2$, it was showed in [BH22, Theorem B] that intermediate von Neumann subalgebras $M\subset\mathscr{N}\subset\mathscr{B}$ are in one-to-one correspondence with intermediate parabolic subgroups $P<Q<G$ via the identification $\mathscr{N}=\operatorname{L}(\Lambda\curvearrowright G/Q)$. In particular, the inclusion $M\subset\mathscr{B}$ retains the abstract poset of all intermediate parabolic subgroups and the rank $\sum_{i=1}^{d}\operatorname{rk}_{k_{i}}(\mathbf{G}_{i})$. However, it is open whether the inclusion $M\subset\mathscr{B}$ retains the Weyl group $\mathscr{W}_{G}$.
Theorem A above implies that the larger inclusion $M\subset\mathscr{M}$ retains the Weyl group $\mathscr{W}_{G}$ via the group isomorphism $\operatorname{Aut}_{M}(\mathscr{M})\cong\mathscr{W}_{G}$ and so in particular the abstract poset of intermediate parabolic subgroups $P<Q<G$. However, Theorem A and [BH22, Theorem B] are of different nature, their proofs do not rely on the same tools and neither theorem implies the other. We emphasize that there is no higher rank assumption in Theorem A while the higher rank assumption is essential in the proof of [BH22, Theorem B] because one uses the noncommutative Nevo–Zimmer theorem [BH19, BBH21] in the simple case ($d=1$) and the dichotomy result for equivariant ucp maps [BBHP20] in the product case ($d\geq 2$).
The present work is also motivated by Connes’ rigidity conjecture for higher rank lattices in semisimple Lie groups.
Connes’ rigidity conjecture.
For every $j\in\{1,2\}$, let $G_{j}$ be a semisimple connected real Lie group with trivial center, no compact factor such that $\operatorname{rk}_{\mathbb{R}}(G_{j})\geq 2$ and let $\Gamma_{j}<G_{j}$ be an irreducible lattice. If $\operatorname{L}(\Gamma_{1})\cong\operatorname{L}(\Gamma_{2})$, then $G_{1}\cong G_{2}$.
For every $j\in\{1,2\}$, let $d_{j}\geq 1$ and for every $i\in\{1,\dots,d_{j}\}$, let $k_{i,j}$ be a local field of characteristic zero, $\mathbf{G}_{i,j}$ a Zariski connected simply connected $k_{i,j}$-isotropic almost $k_{i,j}$-simple algebraic $k_{i,j}$-group and $\mathbf{S}_{i,j}<\mathbf{G}_{i,j}$ a maximal $k_{i,j}$-split torus. Set $G_{j}=\prod_{i=1}^{d_{j}}\mathbf{G}_{i,j}(k_{i,j})$, $H_{j}=\prod_{i=1}^{d_{j}}\mathscr{Z}_{\mathbf{G}_{i,j}}(\mathbf{S}_{i,j})(k_{i,j})$ and $\mathscr{W}_{G_{j}}=\prod_{i=1}^{d_{j}}\mathscr{W}_{\mathbf{G}_{i,j}}$. Let $\Gamma_{j}<G_{j}$ be an irreducible lattice and set $\Lambda_{j}=\Gamma_{j}/\mathscr{Z}(\Gamma_{j})$. Following Theorem A, denote by $\rho_{j}:\mathscr{W}_{G_{j}}\to\operatorname{Aut}_{\operatorname{L}(\Lambda_{j})}(\operatorname{L}(\Lambda_{j}\curvearrowright G_{j}/H_{j}))$ the corresponding group isomorphism. We derive the following consequence of Theorem A which gives new insight towards Connes’ rigidity conjecture.
Corollary B.
Let $\Psi:\operatorname{L}(\Lambda_{1}\curvearrowright G_{1}/H_{1})\to\operatorname{L}(\Lambda_{2}\curvearrowright G_{2}/H_{2})$ be a surjective unital normal $\ast$-isomorphism such that $\Psi(\operatorname{L}(\Lambda_{1}))=\operatorname{L}(\Lambda_{2})$. Then the mapping $\mathscr{W}_{G_{1}}\to\mathscr{W}_{G_{2}}:w\mapsto\rho_{2}^{-1}(\Psi\circ\rho_{1}(w)\circ\Psi^{-1})$ is a group isomorphism.
Let us say a few words about the proof of Theorem A. Let $\Theta\in\operatorname{Aut}_{M}(\mathscr{M})$. In order to prove that there exists $w\in\mathscr{W}_{G}$ such that $\Theta=\Theta_{w}$, it is enough to show that $\Theta(\mathscr{A})=\mathscr{A}$, where $\mathscr{A}=\operatorname{L}^{\infty}(G/H)$. To do this, we proceed in three steps. Firstly, we consider the equivariant normal ucp map $\Phi=\operatorname{E}\circ\Theta|_{\mathscr{A}}:\mathscr{A}\to\mathscr{A}$, where $\operatorname{E}:\mathscr{M}\to\mathscr{A}$ is the canonical faithful normal conditional expectation. By interpreting the main result of [BFGW12] as a rigidity statement for equivariant normal ucp maps, we conclude that $\Phi=\sum_{w\in\mathscr{W}_{G}}\alpha_{w}\theta_{w}$ is a convex combination of the finite group $\left\{\theta_{w}\mid w\in\mathscr{W}_{G}\right\}$. Secondly, we develop a convexity/maximality argument to show that there exist $w\in\mathscr{W}_{G}$ and $u\in\mathscr{U}(\mathscr{M})$ such that $\Theta=\operatorname{Ad}(u)\circ\Theta_{w}$ on $\mathscr{A}$. This second step is purely noncommutative and is reminiscent of the intertwining method in Popa’s deformation/rigidity theory [Po06]. Thirdly, by exploiting the essential freeness of the nonsingular action $\Lambda\times\mathscr{W}_{G}\curvearrowright G/H$, we conclude that $u\in\mathscr{U}(\mathscr{A})$, which finally implies that $\Theta=\Theta_{w}$.
Acknowledgments
CH is grateful to Amine Marrakchi for insightful conversations that led to Proposition 2.7 and to François Thilmany for thought-provoking discussions on the structure and classification of simple algebraic groups. The authors also thank Uri Bader and Itamar Vigdorovich for their valuable comments.
2. Preliminaries
We keep the same notation as in the introduction. For any group action $L\curvearrowright Z$, we denote by $Z^{L}$ the set of all $L$-fixed points in $Z$. For any subgroup $L<R$, the quotient group $\mathscr{N}_{R}(L)/L$ is naturally identified with the group $\operatorname{Aut}_{R}(R/L)$ of all $R$-equivariant bijections of the set $R/L$ via the onto isomorphism
$\mathscr{N}_{R}(L)/L\to\operatorname{Aut}_{R}(R/L):nL\mapsto\left(gL\mapsto gLn^{-1}\right).$
2.1. The Weyl group
In this subsection, we fix a local field $k$ of characteristic zero, $\mathbf{G}$ a Zariski connected simply connected $k$-isotropic almost $k$-simple algebraic $k$-group and $\mathbf{S}<\mathbf{G}$ a maximal $k$-split torus. Set $\mathbf{H}=\mathscr{Z}_{\mathbf{G}}(\mathbf{S})$. The Weyl group of $\mathbf{G}$ relative to $k$ is defined by the formula
$\mathscr{W}_{\mathbf{G}}=\mathscr{N}_{\mathbf{G}}(\mathbf{S})/\mathscr{Z}_{\mathbf{G}}(\mathbf{S}).$
Proposition 2.1.
Keep the same notation as above. The following assertions hold:
  • $(\rm i)$
    We have $\mathscr{N}_{\mathbf{G}}(\mathbf{S})(k)=\mathscr{N}_{\mathbf{G}}(\mathbf{H})(k)$.
  • $(\rm ii)$
    We have the natural identifications
    $\mathscr{W}_{\mathbf{G}}\cong\mathscr{N}_{\mathbf{G}(k)}(\mathbf{S}(k))/\mathbf{H}(k)=\mathscr{N}_{\mathbf{G}(k)}(\mathbf{H}(k))/\mathbf{H}(k).$
    In particular, $\mathscr{W}_{\mathbf{G}}\cong\operatorname{Aut}_{\mathbf{G}(k)}(\mathbf{G}(k)/\mathbf{H}(k))$ coincides the group of all nonsingular $\mathbf{G}(k)$-equivariant automorphisms of $\mathbf{G}(k)/\mathbf{H}(k)$.
  • $(\rm iii)$
    We have $(\mathbf{G}(k)/\mathbf{H}(k))^{\mathbf{H}(k)}=\mathscr{N}_{\mathbf{G}(k)}(\mathbf{H}(k))/\mathbf{H}(k)\cong\mathscr{W}_{\mathbf{G}}$.
Proof.
By [Sh97, Proposition 1.4], every algebraic $k$-group $\mathbf{L}$ admits a maximal (with respect to inclusion) $k$-subgroup $\mathbf{N}$, called the $k$-discompact radical of $\mathbf{L}$, which has no nontrivial $k$-compact quotients. Moreover, $\mathbf{N}<\mathbf{L}$ is a characteristic (hence, normal) subgroup and the quotient $\mathbf{N}/\mathbf{L}$ is $k$-compact.
We claim that $\mathbf{S}$ is the $k$-discompact radical of $\mathbf{H}$. Indeed, denote by $\mathbf{R}$ the $k$-discompact radical of $\mathbf{H}$. Since $\mathbf{S}$ has no nontrivial $k$-compact quotient, it follows that $\mathbf{S}<\mathbf{R}$. Since $\mathbf{S}$ is central in $\mathbf{H}$, we have that $\mathbf{R}/\mathbf{S}$ is a $k$-subgroup of $\mathbf{H}/\mathbf{S}$. Since the characteristic of $k$ is zero, the quotient group $\mathbf{H}(k)/\mathbf{S}(k)$ has finite index in $(\mathbf{H}/\mathbf{S})(k)$. Since $\mathbf{H}(k)/\mathbf{S}(k)$ is compact by [Ma91, Proposition I.2.3.6], it follows that $(\mathbf{H}/\mathbf{S})(k)$ is compact. Then $(\mathbf{R}/\mathbf{S})(k)$ is compact and since $\mathbf{R}$ is the $k$-discompact radical of $\mathbf{H}$, we necessarily have $\mathbf{S}=\mathbf{R}$.
$(\rm i)$ It is plain to see that $\mathscr{N}_{\mathbf{G}}(\mathbf{S})(k)<\mathscr{N}_{\mathbf{G}}(\mathbf{H})(k)$. Conversely, let $n\in\mathscr{N}_{\mathbf{G}}(\mathbf{H})(k)$. Then $\iota_{n}:\mathbf{H}\to\mathbf{H}:h\mapsto nhn^{-1}$ is a $k$-automorphism. Since the $k$-discompact radical $\mathbf{S}$ is a characteristic subgroup of $\mathbf{H}$, it follows that $n\mathbf{S}n^{-1}=\mathbf{S}$ and so $n\in\mathscr{N}_{\mathbf{G}}(\mathbf{S})(k)$.
$(\rm ii)$ Since $\mathbf{S}$ is a Zariski connected $k$-group, $\mathbf{S}(k)$ is Zariski dense in $\mathbf{S}$ by [Ma91, Proposition 1.2.5.3(ii)]. This implies that $\mathscr{N}_{\mathbf{G}(k)}(\mathbf{S}(k))=\mathscr{N}_{\mathbf{G}}(\mathbf{S})(k)$. Likewise, since $\mathbf{H}$ is a Zariski connected $k$-group, we have $\mathscr{N}_{\mathbf{G}(k)}(\mathbf{H}(k))=\mathscr{N}_{\mathbf{G}}(\mathbf{H})(k)$. Using $(\rm i)$, we obtain $\mathscr{N}_{\mathbf{G}(k)}(\mathbf{S}(k))=\mathscr{N}_{\mathbf{G}(k)}(\mathbf{H}(k))$ and so
$\mathscr{N}_{\mathbf{G}(k)}(\mathbf{S}(k))/\mathbf{H}(k)=\mathscr{N}_{\mathbf{G}(k)}(\mathbf{H}(k))/\mathbf{H}(k).$
We know that the image of $\mathscr{N}_{\mathbf{G}}(\mathbf{S})(k)$ under the epimorphism $\mathscr{N}_{\mathbf{G}}(\mathbf{S})\to\mathscr{W}_{\mathbf{G}}$ is all of $\mathscr{W}_{\mathbf{G}}$. This further implies that $\mathscr{N}_{\mathbf{G}}(\mathbf{S})(k)/\mathbf{H}(k)$ coincides with the $\mathscr{N}_{\mathbf{G}}(\mathbf{S})(k)$-orbit of $\mathbf{H}$ in $\mathscr{N}_{\mathbf{G}}(\mathbf{S})/\mathbf{H}$. Therefore, we have the natural identifications
$\mathscr{W}_{\mathbf{G}}\cong\mathscr{N}_{\mathbf{G}(k)}(\mathbf{S}(k))/\mathbf{H}(k)=\mathscr{N}_{\mathbf{G}(k)}(\mathbf{H}(k))/\mathbf{H}(k).$
The moreover part follows from the observation made at the beginning of the section.
$(\rm iii)$ It is plain to see that $\mathscr{N}_{\mathbf{G}(k)}(\mathbf{H}(k))/\mathbf{H}(k)\subset(\mathbf{G}(k)/\mathbf{H}(k))^{\mathbf{H}(k)}$. Conversely, let $g\mathbf{H}(k)\in(\mathbf{G}(k)/\mathbf{H}(k))^{\mathbf{H}(k)}$. Then we have $g^{-1}\mathbf{H}(k)g<\mathbf{H}(k)$. Since $\mathbf{H}$ is a Zariski connected $k$-group, $\mathbf{H}(k)$ is Zariski dense in $\mathbf{H}$ by [Ma91, Proposition 1.2.5.3(ii)] and so $g^{-1}\mathbf{H}g<\mathbf{H}$. By the descending chain condition, we obtain $g^{-1}\mathbf{H}g=\mathbf{H}$ and so $g^{-1}\mathbf{H}(k)g=\mathbf{H}(k)$. Thus, we have $g\mathbf{H}(k)\in\mathscr{N}_{\mathbf{G}(k)}(\mathbf{H}(k))/\mathbf{H}(k)$. ∎
2.2. Algebraic groups and homogeneous spaces
We now come back to the notation we introduced in the introduction. Firstly, we review the work of [BFGW12] regarding the rigidity of group actions on algebraic homogeneous spaces. For every $i\in\{1,\dots,d\}$, let $\mathbf{L}_{i}<\mathbf{G}_{i}$ be a $k_{i}$-subgroup and $\mathbf{V}_{i}$ a $k_{i}$-$\mathbf{G}_{i}$-algebraic variety. Set $L=\prod_{i=1}^{d}\mathbf{L}_{i}(k_{i})$ and $V=\prod_{i=1}^{d}\mathbf{V}_{i}(k_{i})$. Define $\operatorname{Prob}^{0}(V)=V$ and $\operatorname{Prob}^{n+1}(V)=\operatorname{Prob}(\operatorname{Prob}^{n}(V))$ for every $n\in\mathbb{N}$. Then for every $n\in\mathbb{N}$, we may regard $\operatorname{Prob}^{n}(V)$ as a Borel $\Gamma$-space (resp. $G$-space) and we denote by $\operatorname{Map}_{\Gamma}(G/L,\operatorname{Prob}^{n}(V))$ (resp. $\operatorname{Map}_{G}(G/L,\operatorname{Prob}^{n}(V))$) the space of all equivalence classes of $\Gamma$-equivariant (resp. $G$-equivariant) Borel maps $\beta:G/L\to\operatorname{Prob}^{n}(V)$. Assume moreover that $L<G$ is noncompact. Since $\Gamma<G$ is irreducible, Moore’s ergodicity theorem implies that $\Gamma\curvearrowright G/L$ is ergodic (see [HM77]).
Theorem 2.2 ([BFGW12, Theorem 1.6]).
Keep the same notation as above. Then for every $n\in\mathbb{N}$, we have
$\operatorname{Map}_{\Gamma}(G/L,\operatorname{Prob}^{n}(V))=\operatorname{Map}_{G}(G/L,\operatorname{Prob}^{n}(V))\cong\operatorname{Prob}^{n}(V^{L}),$
where $w\in\operatorname{Prob}^{n}(V^{L})$ corresponds to the measurable $G$-equivariant map $\Phi_{w}:G/L\to\operatorname{Prob}^{n}(V):gL\mapsto g\cdot w$.
In particular, in the case when $L=H$, we have
$\operatorname{Aut}_{\Lambda}(G/H)=\operatorname{Aut}_{\Gamma}(G/H)=\operatorname{Aut}_{G}(G/H)\cong\mathscr{W}_{G}.$
Proof.
By assumption, for every $i\in\{1,\dots,d\}$, $\mathbf{G}_{i}$ is a Zariski connected simply connected $k_{i}$-isotropic almost $k_{i}$-simple algebraic $k_{i}$-group. Then [Ma91, Theorem I.1.5.6 and Corollary I.5.6.7] imply that any proper normal subgroup of $\mathbf{G}_{i}(k_{i})$ is contained in $\mathscr{Z}(\mathbf{G}_{i})$ and $\mathbf{G}_{i}(k_{i})$ has no proper finite index subgroup. This further implies that $G$ has no nontrivial algebraic compact factor group ($G$ satisfies condition $(\ast)$ in [BFGW12]) and $G$ has no proper finite index subgroup. By applying [BFGW12, Theorem 1.6] and using the notation therein, we have that $M=G$ and so the conclusion follows.
[…]
$\operatorname{Aut}_{\Lambda}(G/H)=\operatorname{Aut}_{\Gamma}(G/H)=\operatorname{Aut}_{G}(G/H)\cong\mathscr{W}_{G}.$
This finishes the proof. ∎
Since the Weyl group $\mathscr{W}_{G}$ coincides with the group $\operatorname{Aut}_{\Lambda}(G/H)$ of all $\Lambda$-equivariant nonsingular automorphisms of $G/H$, we may consider the well-defined nonsingular action $\Lambda\times\mathscr{W}_{G}\curvearrowright G/H$. The next lemma will be crucial in the proof of our main result.
Lemma 2.3.
The nonsingular action $\Lambda\times\mathscr{W}_{G}\curvearrowright G/H$ is essentially free.
Proof.
Let $(\lambda,w)\in\Lambda\times\mathscr{W}_{G}\setminus\{(e,e)\}$. We need to prove that the fixed-point subset
$\left\{gH\in G/H\mid(\lambda,w)\cdot gH=gH\right\}$
is null in $G/H$. Choose $\gamma=(\gamma_{i})_{i}\in G=\prod_{i=1}^{d}\mathbf{G}_{i}(k_{i})$ so that $\lambda=\gamma\mathscr{Z}(\Gamma)$ and $n=(n_{i})_{i}=\prod_{i=1}^{d}\mathscr{N}_{\mathbf{G}_{i}}(\mathbf{S}_{i})(k_{i})$ so that $w=(n_{i}\mathbf{H}_{i})_{i}$. Since $(\lambda,w)\neq(e,e)$, there exists $i\in\{1,\dots,d\}$ such that $\gamma_{i}\notin\mathscr{Z}(\mathbf{G}_{i})$ or $n_{i}\notin\mathbf{H}_{i}$. Denote by $\pi_{i}:\mathbf{G}_{i}\to\mathbf{G}_{i}/\mathbf{H}_{i}$ the canonical algebraic $k_{i}$-morphism. Set $G_{i}=\mathbf{G}_{i}(k_{i})$, $H_{i}=\mathbf{H}_{i}(k_{i})$ and regard $G_{i}/H_{i}=\mathbf{G}_{i}(k_{i})/\mathbf{H}_{i}(k_{i})=\mathbf{G}_{i}(k_{i})\pi_{i}(e)\subset(\mathbf{G}_{i}/\mathbf{H}_{i})(k_{i})$ as a closed and open subset (see e.g. [Ma91, Proposition I.2.1.4]). Denote by $p_{i}={\pi_{i}}|_{G_{i}}:G_{i}\to G_{i}/H_{i}$ the corresponding restriction map. It suffices to show that the fixed-point subset
$W_{i}=\left\{gH_{i}\in G_{i}/H_{i}\mid\gamma_{i}gH_{i}n_{i}=gH_{i}\right\}$
is null in $G_{i}/H_{i}$. By the general theory of locally compact second countable groups, $W_{i}$ is null in $G_{i}/H_{i}$ if and only if $V_{i}=p_{i}^{-1}(W_{i})$ is null in $G_{i}$.
Denote by $\mathbf{W}_{i}$ the Zariski closure of $W_{i}$ in $\mathbf{G}_{i}/\mathbf{H}_{i}$ and observe that $\mathbf{W}_{i}\subset\left\{g\mathbf{H}_{i}\in\mathbf{G}_{i}/\mathbf{H}_{i}\mid\gamma_{i}g\mathbf{H}_{i}n_{i}=g\mathbf{H}_{i}\right\}$. By [Bo91, Theorem 14.4], $\mathbf{W}_{i}\subset\mathbf{G}_{i}/\mathbf{H}_{i}$ is an algebraic subvariety defined over $k_{i}$ for which $W_{i}\subset G_{i}/H_{i}\cap\mathbf{W}_{i}(k_{i})$. We claim that $\mathbf{W}_{i}$ is a proper subvariety of $\mathbf{G}_{i}/\mathbf{H}_{i}$. Indeed, otherwise if $\mathbf{W}_{i}=\mathbf{G}_{i}/\mathbf{H}_{i}$, then we have
$\mathbf{G}_{i}/\mathbf{H}_{i}=\left\{g\mathbf{H}_{i}\in\mathbf{G}_{i}/\mathbf{H}_{i}\mid\gamma_{i}g\mathbf{H}_{i}n_{i}=g\mathbf{H}_{i}\right\}.$
In particular, we have $n_{i}\mathbf{H}_{i}=\mathbf{H}_{i}n_{i}=\gamma_{i}^{-1}\mathbf{H}_{i}=\mathbf{H}_{i}\gamma_{i}^{-1}$ and so $\gamma_{i}g\mathbf{H}_{i}\gamma_{i}^{-1}=g\mathbf{H}_{i}$ for every $g\mathbf{H}_{i}\in\mathbf{G}_{i}/\mathbf{H}_{i}$. Since $\mathbf{H}_{i}<\mathbf{P}_{i}$, this further implies that $\gamma_{i}\mathbf{P}_{i}\gamma_{i}^{-1}=\mathbf{P}_{i}$. Since $\mathscr{N}_{\mathbf{G}_{i}}(\mathbf{P}_{i})=\mathbf{P}_{i}$, it follows that $\gamma_{i}\in\mathbf{P}_{i}$. Then we have $n_{i}\in\mathscr{N}_{\mathbf{G}_{i}}(\mathbf{S}_{i})\cap\mathbf{P}_{i}=\mathscr{Z}_{\mathbf{G}_{i}}(\mathbf{S}_{i})=\mathbf{H}_{i}$. Moreover, for every $g\mathbf{H}_{i}\in\mathbf{G}_{i}/\mathbf{H}_{i}$, we have $\gamma_{i}g\mathbf{H}_{i}=g\mathbf{H}_{i}$ and so $\gamma_{i}\in\bigcap_{g\in\mathbf{G}_{i}}g\mathbf{H}_{i}g^{-1}$. Since $\mathbf{H}_{i}<\mathbf{G}_{i}$ is a proper $k_{i}$-subgroup, $\bigcap_{g\in\mathbf{G}_{i}}g\mathbf{H}_{i}g^{-1}\lhd\mathbf{G}_{i}$ is a proper normal $k_{i}$-closed subgroup. Since $\mathbf{G}_{i}$ is almost $k_{i}$-simple, we have $\bigcap_{g\in\mathbf{G}_{i}}g\mathbf{H}_{i}g^{-1}=\mathscr{Z}(\mathbf{G}_{i})$ and so $\gamma_{i}\in\mathscr{Z}(\mathbf{G}_{i})$. Therefore, we showed that $\gamma_{i}\in\mathscr{Z}(\mathbf{G}_{i})$ and $n_{i}\in\mathbf{H}_{i}$. This is a contradiction and so $\mathbf{W}_{i}$ is a proper algebraic subvariety of $\mathbf{G}_{i}/\mathbf{H}_{i}$.
[…]
2.3. Normal ucp maps and boundary maps
We recall the following well-known result on equivariant normal ucps maps between abelian von Neumann algebras.
Proposition 2.4.
Let $L$ be a locally compact second countable group and $(X,\nu_{X})$ (resp. $(Y,\nu_{Y})$) a locally compact second countable $L$-space.
To any $L$-equivariant normal ucp map $\Phi:\operatorname{L}^{\infty}(Y,\nu_{Y})\to\operatorname{L}^{\infty}(X,\nu_{X})$ corresponds an essentially unique $L$-equivariant measurable map $\beta:X\to\operatorname{Prob}(Y):x\mapsto\beta_{x}$ such that $\operatorname{Bar}(\beta_{\ast}\nu_{X})=\nu_{X}\circ\Phi\prec\nu_{Y}$ and for $\nu_{X}$-almost every $x\in X$ and every $f\in\operatorname{C}_{0}(Y)$, we have
$\beta_{x}(f)=\Phi(f)(x).$
Proof.
The result is well known when $Y$ is compact. Without loss of generality, we may assume that $Y$ is noncompact and that the topological support of $\nu_{Y}$ is equal to $Y$. Then we have $\operatorname{C}_{0}(Y)\subset\operatorname{L}^{\infty}(Y,\nu_{Y})$. Denote by $Z=Y\sqcup\{\infty\}$ the one-point compactification of $Y$, which is a compact metrizable space. Define the Borel probability measure $\nu_{Z}\in\operatorname{Prob}(Z)$ by $\nu_{Z}|_{Y}=\nu_{Y}$ and $\nu_{Z}(\{\infty\})=0$. Regard $\operatorname{Prob}(Y)\subset\operatorname{Prob}(Z)$ as a Borel subset. Then we have $\operatorname{L}^{\infty}(Y,\nu_{Y})=\operatorname{L}^{\infty}(Z,\nu_{Z})$. Let $\Phi:\operatorname{L}^{\infty}(Y,\nu_{Y})\to\operatorname{L}^{\infty}(X,\nu_{X})$ be an $L$-equivariant normal ucp map. Since the result holds for the compact metrizable space $Z$, there exists an essentially unique $L$-equivariant measurable map $\beta:X\to\operatorname{Prob}(Z):x\mapsto\beta_{x}$ such that $\operatorname{Bar}(\beta_{\ast}\nu_{X})=\nu_{X}\circ\Phi\prec\nu_{Z}$ and for $\nu_{X}$-almost every $x\in X$ and every $f\in\operatorname{C}(Z)$, we have $\beta_{x}(f)=\Phi(f)(x)$. It suffices to prove that $\beta_{x}(\{\infty\})=0$ for $\nu_{X}$-almost every $x\in X$.
We may find a nondecreasing sequence $f_{n}\in\operatorname{C}_{0}(Y)$ of compactly supported functions such that $0\leq f_{n}\leq 1$ for every $n\in\mathbb{N}$ and $f_{n}\to\mathbf{1}_{Y}$ pointwise. For every $n\in\mathbb{N}$, denote by $K_{n}=\operatorname{supp}(f_{n})$ the compact support in $Y$ of $f_{n}\in\operatorname{C}_{0}(Y)$. Lebesgue’s dominated convergence theorem implies that $f_{n}\to\mathbf{1}_{Y}$ strongly in $\operatorname{L}^{\infty}(Y,\nu_{Y})$. Since $\Phi:\operatorname{L}^{\infty}(Y,\nu_{Y})\to\operatorname{L}^{\infty}(X,\nu_{X})$ is normal, we have $\Phi(f_{n})\to\mathbf{1}_{X}$ strongly in $\operatorname{L}^{\infty}(X,\nu_{X})$. Upon taking a subsequence, we may further assume that there is a conull Borel subset $X_{0}\subset X$ for which $\beta_{x}(f_{n})=\Phi(f_{n})(x)\to 1$ for every $x\in X_{0}$. Then for every $x\in X_{0}$, we have $\beta_{x}(f_{n})\leq\beta_{x}(K_{n})\leq 1$ for every $n\in\mathbb{N}$ and so $\lim_{n}\beta_{x}(K_{n})=1$. Since $(K_{n})_{n}$ is nondecreasing, letting $K=\bigcup_{n\in\mathbb{N}}K_{n}\subset Y$, for every $x\in X_{0}$, we have that $\beta_{x}(K)=1$. Thus, for every $x\in X_{0}$, we obtain $\beta_{x}(\{\infty\})\leq\beta_{x}(Z\setminus K)=\beta_{x}(Z)-\beta_{x}(K)=0$. ∎
As a corollary to Theorem 2.2 and Proposition 2.4, we obtain the following rigidity result for equivariant normal ucp maps.
Theorem 2.5.
Let $\Phi:\operatorname{L}^{\infty}(G/H)\to\operatorname{L}^{\infty}(G/H)$ be a $\Lambda$-equivariant normal ucp map. Then there exists a unique tuple $(\alpha_{w})_{w\in\mathscr{W}_{G}}\in[0,1]^{\mathscr{W}_{G}}$ such that $\sum_{w\in\mathscr{W}_{G}}\alpha_{w}=1$ and
$\Phi=\sum_{w\in\mathscr{W}_{G}}\alpha_{w}\theta_{w}.$
Proof.
[…]
2.4. The group measure space construction
We recall the following well-known fact.
Lemma 2.6.
The group $\Lambda$ has infinite conjugacy classes.
Proof.
Let $\gamma\in\Gamma$ be such that its conjugacy class $C(\gamma)=\left\{h\gamma h^{-1}\mid h\in\Gamma\right\}$ is finite. We need to show that $\gamma\in\mathscr{Z}(\Gamma)$. There are two cases to consider.
Firstly, assume that $d=1$. Simply write $\mathbf{G}=\mathbf{G}_{1}$. By assumption on $\mathbf{G}$, since $\mathscr{Z}_{\Gamma}(\gamma)<\Gamma$ has finite index, $\mathscr{Z}_{\Gamma}(\gamma)<G$ is a lattice and so $\mathscr{Z}_{\Gamma}(\gamma)$ is Zariski dense in $\mathbf{G}$ by Borel’s density theorem [Bo60] (see also [Sh97, BDL14]). Since $\mathscr{Z}_{\Gamma}(\gamma)<\mathscr{Z}_{\mathbf{G}}(\gamma)$, this implies that $\mathbf{G}=\mathscr{Z}_{\mathbf{G}}(\gamma)$ and so $\gamma\in\mathscr{Z}(\mathbf{G})$ hence $\gamma\in\mathscr{Z}(\Gamma)$.
Secondly, assume that $d\geq 2$. Let $i\in\{1,\dots,d\}$, set $G_{i}=\mathbf{G}_{i}(k_{i})$ and denote by $q_{i}:G\to G_{i}$ the quotient homomorphism. By assumption, $q_{i}(\Gamma)<G_{i}$ is dense for the analytic topology. Since $\mathscr{Z}_{\Gamma}(\gamma)<\Gamma$ has finite index, it follows that the closure of $q_{i}(\mathscr{Z}_{\Gamma}(\gamma))$ in $G_{i}$ for the analytic topology has finite index in $G_{i}$. Since $G_{i}$ has no proper finite index subgroup (see [Ma91, Corollary I.1.5.7]), $q_{i}(\mathscr{Z}_{\Gamma}(\gamma))<G_{i}$ is dense for the analytic topology. Since $q_{i}(\mathscr{Z}_{\Gamma}(\gamma))<\mathscr{Z}_{G_{i}}(q_{i}(\gamma))$, this implies that $G_{i}=\mathscr{Z}_{G_{i}}(q_{i}(\gamma))$ and so $q_{i}(\gamma)\in\mathscr{Z}(G_{i})$. Since this holds for every $i\in\{1,\dots,d\}$, this further implies that $\gamma\in\mathscr{Z}(G)$ and so $\gamma\in\mathscr{Z}(\Gamma)$. ∎
By assumption, the locally compact group $G$ is unimodular. The closed subgroup $H<G$ is also unimodular because $S$ is central in $H$ and the quotient group $H/S$ is compact. Thus, [BHV08, Corollary B.1.7] implies that the homogeneous space $G/H$ carries a $G$-invariant $\sigma$-finite infinite measure $m$ whose measure class coincides with the unique $G$-invariant measure class on $G/H$. We denote by $\operatorname{E}:\mathscr{M}\to\mathscr{A}:\sum_{\gamma\in\Lambda}a_{\gamma}u_{\gamma}\mapsto a_{e}$ the canonical faithful normal conditional expectation. We may then fix a faithful normal semifinite trace $\operatorname{Tr}$ on $\mathscr{M}$ so that $\operatorname{Tr}\circ\operatorname{E}=\operatorname{Tr}$ and $\operatorname{Tr}|_{\mathscr{A}}=\int_{G/H}\cdot\,{\rm d}m$.
Proposition 2.7.
Keep the same notation as above. Then the inclusion $M\subset\mathscr{M}$ is ergodic in the sense that $M^{\prime}\cap\mathscr{M}=\mathbb{C}1$.
Proof.
Choose a Borel probability measure $\nu\in\operatorname{Prob}(G/H)$ whose measure class coincides with the unique $G$-invariant measure on $G/H$. Regard $\mathscr{M}=\operatorname{L}(\Lambda\curvearrowright G/H)=(\operatorname{L}^{\infty}(G/H)\mathbin{\overline{\otimes}}\operatorname{B}(\ell^{2}(\Lambda)))^{\Lambda}$ as the fixed-point von Neumann subalgebra where the action $\Lambda\curvearrowright\operatorname{L}^{\infty}(G/H)\mathbin{\overline{\otimes}}\operatorname{B}(\ell^{2}(\Lambda))$ is defined by the formula
$\forall\gamma\in\Lambda,\forall F\in\operatorname{L}^{\infty}(G/H)\mathbin{\overline{\otimes}}\operatorname{B}(\ell^{2}(\Lambda)),\quad\gamma\cdot F=\left(\sigma_{\gamma}\otimes\operatorname{Ad}(Ju_{\gamma}J)\right)(F).$
Then we may regard $M^{\prime}\cap\mathscr{M}=\operatorname{L}(\Lambda)^{\prime}\cap\operatorname{L}(\Lambda\curvearrowright G/H)$ as the von Neumann algebra of all equivalence classes of $\Lambda$-equivariant measurable functions $F:G/H\to J\operatorname{L}(\Lambda)J$, where we consider the action $\Lambda\curvearrowright J\operatorname{L}(\Lambda)J$ given by $\rho_{\gamma}=\text{Ad}(Ju_{\gamma}J)$. If we endow $J\operatorname{L}(\Lambda)J$ with the $\|\cdot\|_{2}$-norm, then $J\operatorname{L}(\Lambda)J$ is separable and $\rho$ is an isometric action.
Since $G=\prod_{i=1}^{d}\mathbf{G}_{i}(k_{i})$ and $H=\prod_{i=1}^{d}\mathbf{H}_{i}(k_{i})$, where $\mathbf{G}_{1},\dots,\mathbf{G}_{d}$ are chosen as in the introduction, and since $\Gamma<G$ is a lattice, [BG14, Theorem 6.6 and Corollary 6.7] imply that the nonsingular action $\Gamma\curvearrowright G/H$ is metrically ergodic, in the sense of [BG14, Definition 6.5], and so is $\Lambda\curvearrowright G/H$. This shows that any $\Lambda$-equivariant measurable function $F:G/H\to J\operatorname{L}(\Lambda)J$ is essentially constant. Since $\Lambda$ has infinite conjugacy classes by Lemma 2.6, this further implies that $M^{\prime}\cap\mathscr{M}=\mathbb{C}1$. ∎
3. Proofs of the main results
Let $\mathscr{N}$ be a von Neumann algebra and denote by $\mathscr{V}(\mathscr{N})$ the set of partial isometries of $\mathscr{N}$. Endow the set $\mathscr{V}(\mathscr{N})$ with the following order relation: for all $u_{1},u_{2}\in\mathscr{V}(\mathscr{N})$, we have $u_{1}\leq u_{2}$ if and only if $u_{1}=u_{2}u_{1}^{*}u_{1}$. It is well known that the partially ordered set $(\mathscr{V}(\mathscr{N}),\leq)$ is inductive.
3.1. Proof of Theorem A
We still denote by $\sigma:\Lambda\curvearrowright\mathscr{M}$ the conjugation action that naturally extends the action $\Lambda\curvearrowright\mathscr{A}$.
Let $\Theta\in\operatorname{Aut}_{M}(\mathscr{M})$. Then $\Phi=\operatorname{E}\circ\Theta|_{\mathscr{A}}:\mathscr{A}\to\mathscr{A}$ is a $\Lambda$-equivariant normal ucp map. By Theorem 2.5, there exists a unique tuple $(\alpha_{w})_{w\in\mathscr{W}_{G}}\in[0,1]^{\mathscr{W}_{G}}$ such that $\sum_{w\in\mathscr{W}_{G}}\alpha_{w}=1$ and $\Phi=\sum_{w\in\mathscr{W}_{G}}\alpha_{w}\theta_{w}$. Let $w_{0}\in\mathscr{W}_{G}$ be such that $\alpha_{w_{0}}>0$. Upon replacing $\Theta$ by $\Theta\circ\Theta_{w_{0}}^{-1}$, without loss of generality, we may further assume that $\alpha_{e}>0$. We need to show that $\Theta=\operatorname{id}_{\mathscr{M}}$.
Firstly, we use a convexity argument to prove that there exists a partial isometry $v\in\mathscr{V}(\mathscr{M})$ such that $\Theta(a)v=va$ for every $a\in\mathscr{A}$. Indeed, let $p\in\mathscr{A}$ be a nonzero finite trace projection and denote by $\mathscr{C}$ the weak closure of the convex hull of the set
$\mathscr{S}=\left\{\left(\Theta(u)pu^{*},(\Theta_{w}(u)pu^{*})_{w\in\mathscr{W}_{G}\setminus\{e\}}\right)\mid u\in\mathscr{U}(\mathscr{A})\right\}\subset\mathscr{M}^{\oplus n},$
where $n=|\mathscr{W}_{G}|$. Denote by $\tau=\operatorname{Tr}^{\oplus n}$ the canonical faithful normal semifinite trace on $\mathscr{M}^{\oplus n}$. Since $\mathscr{S}$ is both uniformly bounded and $\|\cdot\|_{2,\tau}$-bounded, it follows that the image of $\mathscr{C}$ in the Hilbert space $\operatorname{L}^{2}(\mathscr{M}^{\oplus n},\tau)$ is closed. Denote by $(c,(c_{w})_{w\in\mathscr{W}_{G}\setminus\{e\}})$ the unique element of minimal $\|\cdot\|_{2,\tau}$-norm. Since for every $u\in\mathscr{U}(\mathscr{A})$, we have
$\left\|\left(\Theta(u)cu^{*},(\Theta_{w}(u)c_{w}u^{*})_{w\in\mathscr{W}_{G}\setminus\{e\}}\right)\right\|_{2,\tau}=\left\|\left(c,(c_{w})_{w\in\mathscr{W}_{G}\setminus\{e\}}\right)\right\|_{2,\tau}$
it follows that
$\left(\Theta(u)cu^{*},(\Theta_{w}(u)c_{w}u^{*})_{w\in\mathscr{W}_{G}\setminus\{e\}}\right)=\left(c,(c_{w})_{w\in\mathscr{W}_{G}\setminus\{e\}}\right).$
Let $w\in\mathscr{W}_{G}\setminus\{e\}$. Then for every $u\in\mathscr{U}(\mathscr{A})$, we have $\theta_{w}(u)c_{w}=\Theta_{w}(u)c_{w}=c_{w}u$. Since the action $\mathscr{W}_{G}\curvearrowright G/H$ is essentially free (see Lemma 2.3), it follows that $c_{w}=0$. For every $u\in\mathscr{U}(\mathscr{A})$, we have
$\operatorname{E}(\Theta(u)pu^{*})=\operatorname{E}(\Theta(u))pu^{*}=\alpha_{e}p+\sum_{w\in\mathscr{W}_{G}\setminus\{e\}}\Theta_{w}(u)pu^{*}.$
Using convexity and taking limits and since $c_{w}=0$ for every $w\in\mathscr{W}_{G}\setminus\{e\}$, we obtain $\operatorname{E}(c)=\alpha_{e}p\neq 0$ and so $c\neq 0$. Write $c=v|c|$ for the polar decomposition of $c\in\mathscr{M}$. Then $v\neq 0$ and $\Theta(u)v=vu$ for every $u\in\mathscr{U}(\mathscr{A})$. Therefore, we have $\Theta(a)v=va$ for every $a\in\mathscr{A}$. Since $\mathscr{A}^{\prime}\cap\mathscr{M}=\mathscr{A}$, it follows that $v^{*}v\in\mathscr{A}^{\prime}\cap\mathscr{M}=\mathscr{A}$ and $vv^{*}\in\Theta(\mathscr{A})^{\prime}\cap\mathscr{M}=\Theta(\mathscr{A})$. Moreover, we have $\Theta(v^{*}v)v=v$ which implies that $\Theta(v^{*}v)vv^{*}=vv^{*}$ and so $vv^{*}\leq\Theta(v^{*}v)$. Let $r\in\mathscr{A}$ be the unique projection such that $vv^{*}=\Theta(r)$. Then we have $\Theta(r)v=vr$ which implies that $v^{*}v=v^{*}vr$ and so $v^{*}v\leq r$. This shows that $\Theta(v^{*}v)\leq\Theta(r)=vv^{*}$ and so $\Theta(v^{*}v)=vv^{*}$. Upon cutting down $v$ on the right hand side by a nonzero finite trace projection in $\mathscr{A}$, we may further assume that $v\neq 0$ and $\operatorname{Tr}(v^{*}v)=\operatorname{Tr}(vv^{*})<+\infty$. Since $\Theta|_{\operatorname{L}(\Lambda)}=\operatorname{id}_{\operatorname{L}(\Lambda)}$, we have that $\Theta(a)\sigma_{\gamma}(v)=\sigma_{\gamma}(v)a$ for every $a\in\mathscr{A}$ and every $\gamma\in\Lambda$.
Secondly, we glue together the partial isometries $\sigma_{\gamma}(v)$ for $\gamma\in\Lambda$ in order to construct a unitary $u\in\mathscr{U}(\mathscr{M})$ such that $\Theta(a)=uau^{*}$ for every $a\in\mathscr{A}$. Indeed, fix an enumeration $\Lambda=\left\{\gamma_{n}\mid n\in\mathbb{N}\right\}$ such that $\gamma_{0}=e$. For every $n\in\mathbb{N}$, set $p_{n}=\bigvee_{0\leq k\leq n}\sigma_{\gamma_{k}}(v^{*}v)\in\mathscr{A}$, $q_{n}=\bigvee_{0\leq k\leq n}\sigma_{\gamma_{k}}(vv^{*})\in\Theta(\mathscr{A})$ and observe that $\max(\operatorname{Tr}(p_{n}),\operatorname{Tr}(q_{n}))<+\infty$. In particular, we have $p_{n}\neq 1$ and $q_{n}\neq 1$. By induction over $n\in\mathbb{N}$, we construct a nondecreasing sequence of partial isometries $(u_{n})_{n}$ in $\mathscr{V}(\mathscr{M})$ such that $\Theta(a)u_{n}=u_{n}a$ for every $a\in\mathscr{A}$, $u_{n}^{*}u_{n}=p_{n}$, $u_{n}u_{n}^{*}=q_{n}$. Set $u_{0}=v$. Assume that $u_{n}\in\mathscr{V}(\mathscr{M})$ has been constructed. The same reasoning as in the previous paragraph shows that $\Theta(p_{n})=\Theta(u_{n}^{*}u_{n})=u_{n}u_{n}^{*}=q_{n}$. Set $r_{n}=p_{n+1}-p_{n}=\sigma_{\gamma_{n+1}}(v^{*}v)-\sigma_{\gamma_{n+1}}(v^{*}v)p_{n}\in\mathscr{A}$ and $s_{n}=q_{n+1}-q_{n}=\sigma_{\gamma_{n+1}}(vv^{*})-\sigma_{\gamma_{n+1}}(vv^{*})q_{n}\in\Theta(\mathscr{A})$. We have
$\Theta(r_{n})$ $=\sigma_{\gamma_{n+1}}(\Theta(v^{*}v))-\sigma_{\gamma_{n+1}}(\Theta(v^{*}v))\Theta(p_{n})$
$=\sigma_{\gamma_{n+1}}(vv^{*})-\sigma_{\gamma_{n+1}}(vv^{*})q_{n}$
$=s_{n}.$
We define the partial isometry $u_{n+1}\in\mathscr{V}(\mathscr{M})$ by
$q_{n}u_{n+1}p_{n}=u_{n}\quad\text{and}\quad s_{n}u_{n+1}r_{n}=s_{n}\sigma_{\gamma_{n+1}}(v)r_{n}.$
We have
$(s_{n}\sigma_{\gamma_{n+1}}(v)r_{n})^{*}(s_{n}\sigma_{\gamma_{n+1}}(v)r_{n})$ $=r_{n}\sigma_{\gamma_{n+1}}(v)^{*}s_{n}\sigma_{\gamma_{n+1}}(v)r_{n}$
$=r_{n}\sigma_{\gamma_{n+1}}(v)^{*}\Theta(r_{n})\sigma_{\gamma_{n+1}}(v)r_{n}$
$=r_{n}\sigma_{\gamma_{n+1}}(v^{*}v)r_{n}$
$=r_{n}.$
Likewise, we have
$(s_{n}\sigma_{\gamma_{n+1}}(v)r_{n})(s_{n}\sigma_{\gamma_{n+1}}(v)r_{n})^{*}=s_{n}.$
Then it is plain to see that $\Theta(a)u_{n+1}=u_{n+1}a$ for every $a\in\mathscr{A}$, $u_{n+1}^{*}u_{n+1}=p_{n}+r_{n}=p_{n+1}$ and $u_{n+1}u_{n+1}^{*}=q_{n}+s_{n}=q_{n+1}$. By construction, we also have $u_{n}\leq u_{n+1}$. Therefore, by induction, we have constructed the desired nondecreasing sequence $(u_{n})_{n}$ in $\mathscr{V}(\mathscr{M})$. Then we may define $u=\sup u_{n}\in\mathscr{V}(\mathscr{M})$ and observe that $u^{*}u=\bigvee_{\gamma\in\Lambda}\sigma_{\gamma}(v^{*}v)$ and $uu^{*}=\bigvee_{\gamma\in\Lambda}\sigma_{\gamma}(vv^{*})$. Since the nonsingular action $\Lambda\curvearrowright G/H$ is ergodic, it follows that $u^{*}u=1=uu^{*}$ and so $u\in\mathscr{U}(\mathscr{M})$. Then we have $\Theta(a)=uau^{*}$ for every $a\in\mathscr{A}$.
Thirdly, we exploit the fact that the nonsingular action $\Lambda\times\mathscr{W}_{G}\curvearrowright G/H$ is essentially free to show that $u\in\mathscr{U}(\mathscr{A})$ and $\Theta=\operatorname{id}_{\mathscr{M}}$. Indeed, write $u=\sum_{\gamma\in\Lambda}a_{\gamma}u_{\gamma}$ for the Fourier expansion of $u\in\mathscr{A}\rtimes\Lambda=\mathscr{M}$. Then for every $a\in\mathscr{A}$, we have
$\sum_{w\in\mathscr{W}_{G}}\alpha_{w}\theta_{w}(a)=\operatorname{E}(\Theta(a))=\operatorname{E}(uau^{*})=\sum_{\gamma\in\Gamma}|a_{\gamma}|^{2}\sigma_{\gamma}(a).$
Applying Proposition 2.4, denote by $\beta:G/H\to\operatorname{Prob}(G/H):b\mapsto\beta_{b}$ the essentially unique measurable map corresponding to the normal ucp map $\Phi=\operatorname{E}\circ\Theta|_{\mathscr{A}}:\mathscr{A}\to\mathscr{A}$. Then for $\nu$-almost every $b\in G/H$, we have
$\sum_{\gamma\in\Gamma}|a_{\gamma}|^{2}(b)\,\delta_{\gamma^{-1}b}=\beta_{b}=\sum_{w\in\mathscr{W}_{G}}\alpha_{w}\,\delta_{w^{-1}b}.$
For $\nu$-almost every $b\in G/H$, we consider the set of atoms of the probability measure $\beta_{b}$. Then Lemma 2.3 further implies that for every $w\neq e$, we have $\alpha_{w}=0$ and every $\gamma\neq e$, we have $a_{\gamma}=0$ $\nu$-almost everywhere. Moreover, we have $\alpha_{e}=1=|a_{e}|^{2}$ and so $u=a_{e}\in\mathscr{U}(\mathscr{A})$. Thus, $\Theta|_{\mathscr{A}}=\operatorname{id}_{\mathscr{A}}$ and so $\Theta=\operatorname{id}_{\mathscr{M}}$. This finishes the proof of Theorem A.
3.2. Proof of Corollary B
Let $\Psi:\operatorname{L}(\Lambda_{1}\curvearrowright G_{1}/H_{1})\to\operatorname{L}(\Lambda_{2}\curvearrowright G_{2}/H_{2})$ be a surjective unital normal $\ast$-isomorphism such that $\Psi(\operatorname{L}(\Lambda_{1}))=\operatorname{L}(\Lambda_{2})$. For every $j\in\{1,2\}$, set $M_{j}=\operatorname{L}(\Lambda_{j})\subset\operatorname{L}(\Lambda_{j}\curvearrowright G_{j}/H_{j})=\mathscr{M}_{j}$. The map
$\operatorname{Aut}_{M_{1}}(\mathscr{M}_{1})\to\operatorname{Aut}_{M_{2}}(\mathscr{M}_{2}):\Theta\mapsto\Psi\circ\Theta\circ\Psi^{-1}$
is a group isomorphism. By Theorem A, the mapping $\mathscr{W}_{G_{1}}\to\mathscr{W}_{G_{2}}:w\mapsto\rho_{2}^{-1}(\Psi\circ\rho_{1}(w)\circ\Psi^{-1})$ is a group isomorphism.
3.3. Further rigidity results
In this subsection, we record a von Neumann algebraic rigidity result regarding isomorphisms between group measure von Neumann algebras arising from actions of irreducible lattices on algebraic homogeneous spaces.
Keep the same notation as before. For every $j\in\{1,2\}$ and every $i\in\{1,\dots,d\}$, let $\mathscr{Z}(\mathbf{G}_{i})<\mathbf{L}_{i,j}<\mathbf{G}_{i}$ be a proper Zariski connected $k_{i}$-subgroup. Set $L_{j}=\prod_{i=1}^{d}\mathbf{L}_{i,j}(k_{i})<G$. Since $\mathscr{Z}(\Gamma)<L_{j}$, we may consider the nonsingular action $\Lambda\curvearrowright G/L_{j}$. Assume moreover that $L_{j}<G$ is noncompact. Since $\Gamma<G$ is irreducible, Moore’s ergodicity theorem implies that $\Lambda\curvearrowright G/L_{j}$ is ergodic (see [HM77]).
Theorem 3.1.
Let $\Psi:\operatorname{L}(\Lambda\curvearrowright G/L_{1})\to\operatorname{L}(\Lambda\curvearrowright G/L_{2})$ be a surjective unital normal $\ast$-isomorphism such that $\Psi|_{\operatorname{L}(\Lambda)}=\operatorname{id}_{\operatorname{L}(\Lambda)}$. Then there exists $g\in G$ such that $gL_{1}g^{-1}=L_{2}$.
Proof.
For every $j\in\{1,2\}$, we may choose a Borel probability measure $\nu_{j}\in\operatorname{Prob}(G/L_{j})$ whose measure class coincides with the unique $G$-invariant measure class on $G/L_{j}$. Denote by $\operatorname{E}_{j}:\operatorname{L}(\Lambda\curvearrowright G/L_{j})\to\operatorname{L}^{\infty}(G/L_{j})$ the canonical $\Lambda$-equivariant faithful normal conditional expectation. Consider the $\Lambda$-equivariant normal ucp map $\Phi=\operatorname{E}_{2}\circ\Psi|_{\operatorname{L}^{\infty}(G/L_{1})}:\operatorname{L}^{\infty}(G/L_{1})\to\operatorname{L}^{\infty}(G/L_{2})$. By Proposition 2.4, there exists an essentially unique $\Lambda$-equivariant measurable map $\beta:G/L_{2}\to\operatorname{Prob}(G/L_{1}):b\mapsto\beta_{b}$ such that $\operatorname{Bar}(\beta_{\ast}\nu_{2})=\nu_{2}\circ\Phi\prec\nu_{1}$ and for $\nu_{2}$-almost every $b\in G/L_{2}$ and every $f\in\operatorname{C}_{0}(G/L_{1})$, we have $\beta_{b}(f)=\Phi(f)(b)$. By Theorem 2.2, we have $(G/L_{1})^{L_{2}}\neq\emptyset$ and so there exists $g\in G$ such that $g^{-1}L_{2}g<L_{1}$. Likewise, by considering the $\Lambda$-equivariant normal ucp map $\operatorname{E}_{1}\circ\Psi^{-1}|_{\operatorname{L}^{\infty}(G/L_{2})}:\operatorname{L}^{\infty}(G/L_{2})\to\operatorname{L}^{\infty}(G/L_{1})$, there exists $h\in G$ such that $h^{-1}L_{1}h<L_{2}$. Set $x=gh\in G$ and observe that $x^{-1}L_{2}x<L_{2}$. Write $x=(x_{i})_{i}\in\prod_{i=1}^{d}\mathbf{G}_{i}(k_{i})$. For every $i\in\{1,\dots,d\}$, since $\mathbf{L}_{i,2}(k_{i})$ is Zariski dense in $\mathbf{L}_{i,2}$ by [Ma91, Proposition I.2.5.3], we have $x_{i}^{-1}\mathbf{L}_{i,2}x_{i}<\mathbf{L}_{i,2}$. By the descending chain condition, we have $x_{i}^{-1}\mathbf{L}_{i,2}x_{i}=\mathbf{L}_{i,2}$ and so $x_{i}^{-1}\mathbf{L}_{i,2}(k_{i})x_{i}=\mathbf{L}_{i,2}(k_{i})$. This further implies that $x^{-1}L_{2}x=L_{2}$ and so $g^{-1}L_{2}g=L_{1}$. ∎
Let us point out that the assumption that $\Psi|_{\operatorname{L}(\Lambda)}=\operatorname{id}_{\operatorname{L}(\Lambda)}$ cannot be dropped in general. Indeed, when $L_{1}$ and $L_{2}$ are amenable and unimodular, the group measure space von Neumann algebras $\operatorname{L}(\Lambda\curvearrowright G/L_{1})$ and $\operatorname{L}(\Lambda\curvearrowright G/L_{2})$ are both $\ast$-isomorphic to the unique AFD type $\rm II_{\infty}$} factor.
Remark 3.2.
We point out that Theorem A holds more generally when for every $i\in\{1,\dots,d\}$, $\mathbf{H}_{i}=\mathscr{Z}_{\mathbf{G}_{i}}(\mathbf{S}_{i})<\mathbf{G}_{i}$ is replaced by any Zariski connected $k_{i}$-subgroup $\mathbf{L}_{i}<\mathbf{G}_{i}$ for which $\mathscr{N}_{\mathbf{G}_{i}}(\mathbf{L}_{i})/\mathbf{L}_{i}$ is finite, every coset of $\mathscr{N}_{\mathbf{G}_{i}}(\mathbf{L}_{i})/\mathbf{L}_{i}$ is represented by an element rational over $k_{i}$ and $L=\prod_{i=1}^{d}\mathbf{L}_{i}(k_{i})$ is noncompact.